OpenStudy (anonymous):

Solve the system of equations. 2x + 3y = 1 y = 3x + 15 ( , )

OpenStudy (anonymous):

@jyar @nincompoop @Preetha @Kainui @ayyookyndall @butterflydreamer @Callisto @iGreen @Bananas1234 @MARC_ @kohai

OpenStudy (anonymous):

help?

OpenStudy (haseeb96):

2x+3y=1 .......Equation #1 y=3x+15.......equation #2 now put y=3x+15 into Equation #1 at the place of y 2x +3(3x+15)=1 2x +9x +45 =1 11x+45-1=0 11x+44=0 11x=-44 x=-44/11 x=-4 now again put x=-4 into equation #2 at the place of y y=3(-4) +15 y=-12+15 y=3

OpenStudy (haseeb96):



OpenStudy (anonymous):

Thank you @Haseeb96 :D

OpenStudy (haseeb96):

Welcome my brother :)

OpenStudy (cwrw238):

@Haseeb96 - you have answered the question in detail but it would be better if you encouraged/helped Bilaal.K to do some of it himself

OpenStudy (cwrw238):

It's by practicing that you really learn math